# Learning Dynamics of Meta-Learning in Small Model Pretraining

### v1

@@ 6 Learning Dynamics
The small and medium models show clear signs of effective meta-learning. Support accuracy gradually increases and stabilizes around 6–7%, while query accuracy climbs steadily above 40%. This pattern indicates that the models are internalizing a useful task prior, and show smooth convergence with relatively little instability.
The tiny model displays a distinct failure mode. While its support accuracy rises modestly, its query accuracy remains stagnant, hovering just above chance (10%). This suggests the model memorizes support examples but fails to learn task-generalizable features-a canonical symptom of underparameterization in meta-learning (Finn et al., 2017; Rajeswaran et al., 2019). In effect, it lacks the representational bandwidth to encode a shared inductive bias across tasks.
The large model exhibits more complex dynamics. Although it achieves high query accuracy-eventually surpassing 50%-its learning curve is noisier, with sharper fluctuations in both support and query accuracy. These instabilities may arise from the interaction between large-scale parameter updates and stochastic task sampling. While still effective overall, this suggests that large models may require additional stabilization strategies during meta-training (e.g., adaptive inner-loop learning rates or better task normalization).
Interestingly, the large model’s query accuracy (Figure 4) reveals a distinct late-phase jump after 4,500 steps-a grokking-like effect where generalization rapidly improves after a prolonged plateau. This turning point coincides exactly with a stabilization in the head weight variance, suggesting that the model eventually consolidates a useful episodic prior. Such behavior echoes findings in grokking literature (Power et al., 2022; Nanda et al., 2023), where test performance lags training loss for an extended period before suddenly aligning. In the MAML setting, this may correspond to the model first learning how to adapt, before learning to generalize from adaptation.
Taken together, these patterns confirm that meta-learning is most stable within a mid-capacity regime. Models must be large enough to encode reusable structure, but not so large that their learning becomes erratic. These insights help contextualize downstream findings: the best generalization often arises from models that strike a balance between representational power and stable task-level adaptation.
@@
To better understand how episodic adaptation pressures model structure, we track the evolution of classifier head weights across meta-training (Figure 5). Since the head is re-initialized and adapted in every episode, its long-term statistics reflect how the outer loop consolidates across-task regularities.
The top panel shows the standard deviation of head weights. All models exhibit growth in weight variance, indicating increasing expressivity in the task-specific head. The small model diverges most sharply, with its weight variance surpassing all others after 2k steps. This suggests an over-specialization effect: the model learns to adapt aggressively to each task, potentially at the cost of stability.
In the lower panel, the mean of the head weights remains near zero for most models, but the tiny model is an outlier. It accumulates a strong bias in one direction over training, indicating that its head converges toward a fixed mapping that is minimally updated across episodes. This aligns with earlier diagnostics showing that its gradient norms collapse early in training.
These dynamics reinforce the idea that episodic MAML induces a scale-sensitive tradeoff: in higher-capacity models, episodic gradients drive generalizable structure into the shared initialization; in lower-capacity models, this same pressure can cause drift or collapse.
@@
Across all MAML-pretrained models, a decline in PER over training is observed, indicating that attention representations become progressively lower-rank. But in the large model, this collapse is not smooth. Instead, there is an abrupt, synchronized drop in both PER and Paloma perplexity at roughly step 3000, which follows an earlier rise in perplexity, and yields a “double descent" Belkin et al. (2019) shape (Figure 2). This kind of dynamic-initial generalization, degradation, and then sharp re-convergence-mirrors phenomena seen in grokking and mechanistic phase change literature (Nanda et al., 2023; Power et al., 2022).
We interpret this behavior as a representational phase transition: the model initially fits the objective using diffuse, high-dimensional representations, which are later compressed into task-specialized, low-rank structures. The descent in PER lags behind the initial perplexity gains, and only after this drop does the second descent in Paloma begin. There is no strong evidence of a comparable phase transition in the vanilla models. While the large and medium variants show mild inflection points in loss and perplexity around step 3000, these are gradual and lack the coordinated sharpness seen in the MAML-trained models.
This suggests that MAML’s bilevel updates and episodic task pressure may help reorganize the optimization landscape to favor discrete qualitative shifts in representation. As explored in Olsson et al. (2022); Wang et al. (2024); Hoogland et al. (2024), model training often proceeds in qualitatively distinct stages: from brute-force fitting, to intermediate rule memorization, to compressed algorithmic abstraction. The drop in PER may signal such a transition—from early diffuse representations to compressed heads tuned to solve the repeated structure of SMLMT episodes. This representational transition is also reflected in the model’s adaptation performance. Around the same step where PER and Paloma perplexity undergo a sharp drop (step 3000), both support and query accuracies rise abruptly (see Figure 4). Prior to this point, query accuracy remains relatively flat, indicating that the model struggles to generalize from support to query examples. But after the phase transition, the model rapidly learns to extrapolate, with query accuracy climbing from near random to over 0.5.
This synchrony across metrics—loss, perplexity, support/query accuracy, and PER—provides compelling evidence of a coordinated phase shift in the model’s learning trajectory. When looking into more granular checkpoints (Figure 6), there is clearer evidence that the model transitions from an early stage where it relies on diffuse representations to a later stage where it reorganizes both its representations and update paths into a lower-dimensional, more modular form capable of few-shot generalization. That said, this phase behavior appears scale-sensitive as it is absent in smaller scales. This suggests that the capacity to reorganize may be gated by scale, and that below a certain threshold, the inductive pressure of MAML induces collapse rather than modularization.
@@ 7 Conclusion
This study set out to determine whether first–order MAML can make the pretraining of sub-billion–parameter decoder LMs both faster and more intelligible. The evidence gathered across four Pico scales indicates that it can. When the hybrid meta-objective is interleaved with ordinary next-token prediction, every model from 11M to 570M parameters reaches the same cross-entropy loss noticeably sooner than its vanilla counterpart, and the larger two variants carry a two-to-three–point F1 advantage into Universal NER. Further, it shows gains in transfer to low-resource languages, which has the potential to improve the equitability of language technology. Equally important, the spectral logs reveal a striking "diversify-then-compress" shift part-way through training. The moment at which the effective-rank knee appears turns out to be a reliable signal of the final NER score, providing a window into the model’s developing inductive structure that is absent from ordinary loss curves.
Several natural extensions suggest themselves. A first step is to learn whether the same phase transition re-emerges when the corpus is multilingual, which would clarify why cross-script transfer remains the weak point of the present models. Varying which backbone layers adapt, how many steps they receive and how frequently episodes are interleaved may unlock better compute–capability trade-offs. Finally, the clear correlation between the effective-rank collapse and downstream utility hints that spectral diagnostics might serve as a self-supervised early-stopping signal.
@@ Acknowledgments
This work was supported by a grant from the Accelerate Programme for Scientific Discovery, made possible by a donation from Schmidt Futures. David Demitri Africa is supported by the Cambridge Trust and the Jardine Foundation. Richard Diehl Martinez is supported by the Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation).
@@ Appendix A Pseudocode
Distributed Subset Masked Language Modeling Tasks (SMLMT) Training
Distributed Autoregressive (AR) Training
@@ C.1 Supplementary Tables
C.2 Pretraining Results
We present the unedited pretraining indicators for each pico-maml-decoder model below, as logged on WandB.
C.3 Downstream Evaluation
I present the full downstream evaluation results below, ordered by fine-tuning regime, pretraining setup, and model size.
C.4 Learning Dynamics
We present the learning dynamics indicators for each pico-maml-decoder model below, as logged on WandB.
Appendix D Default pico-maml-train Configurations
| Category | Parameter | Default Value |
| Model | Model Type | pico_decoder |
| Hidden Dimension () | 768 | |
| Number of Layers () | 12 | |
| Vocabulary Size | 50,304 | |
| Sequence Length | 2,048 | |
| Attention Heads | 12 | |
| Key/Value Heads | 4 | |
| Activation Hidden Dim | 3,072 | |
| Normalization Epsilon | ||
| Positional Embedding Theta | 10,000.0 | |
| Training | Optimizer | AdamW |
| Learning Rate | ||
| LR Scheduler | Linear w/ Warmup | |
| Warmup Steps | 2,500 | |
| Gradient Accumulation Steps | 128 | |
| Max Training Steps | 200,000 | |
| Precision | BF16 Mixed | |
| Data | Dataset Name | pico-lm/pretokenized-dolma |
| Batch Size | 1,024 | |
| Tokenizer | allenai/OLMo-7B-0724-hf | |
| Checkpointing | Auto Resume | True |
| Save Every N Steps | 100 | |
| Learning Dynamics Layers |
"attention.v_proj",
"attention.o_proj", "swiglu.w_2" |
| Learning Dynamics Eval Data | pico-lm/pretokenized-paloma-tinsy | |
| Evaluation | Metrics | ["paloma"] |
| Paloma Dataset Name | pico-lm/pretokenized-paloma-tinsy | |
| Eval Batch Size | 16 | |
| Monitoring | Logging Level | INFO |
| Log Every N Steps | 100 | |
| Meta-Learning | Enabled | True |
| Hybrid Ratio | 0.5 | |
| Inner Steps () | 10 | |
| Inner Learning Rate | 0.001 | |
| Support Shots () | 4 | |
| Query Ways () | 32 | |
| Classifier Head Layers | 4 | |
| Classifier Head Hidden Dim | 128 | |
| Classifier Head Dropout | 0.1 | |
| Classifier Head Init Method | xavier | |
| Monitoring | Logging Level | INFO |
| Log Every N Steps | 100 |
Appendix E pico-maml-decoder Models Comparison
| Pico-MAML-Decoder Model Comparison | ||||
| Attribute | tiny | small | medium | large |
| Parameter Count | 11M | 65M | 181M | 570M |
| Hidden Dimension () | 96 | 384 | 768 | 1536 |
| Feed-forward Dim | 384 | 1536 | 3072 | 6144 |
| Training Time (6k steps) | 10h | 15h | 16h | 25h |

### v2

@@ 6 Learning Dynamics
The small and medium models show clear signs of effective meta-learning. Support accuracy gradually increases and stabilizes around 6–7%, while query accuracy climbs steadily above 40%. This pattern indicates that the models are internalizing a useful task prior, and show smooth convergence with relatively little instability.The tiny model displays a distinct failure mode. While its support accuracy rises modestly, its query accuracy remains stagnant, hovering just above chance (10%). This suggests the model memorizes support examples but fails to learn task-generalizable features-a canonical symptom of underparameterization in meta-learning (Finn et al., 2017; Rajeswaran et al., 2019). In effect, it lacks the representational bandwidth to encode a shared inductive bias across tasks.The large model shows a late-phase rise in query accuracy after 4,500 steps, coinciding with stabilization of head-weight variance. This suggests a phase-like reorganization where the model consolidates a useful episodic prior after a prolonged plateau. In the MAML setting, this may correspond to the model first learning how to adapt, before learning to generalize from adaptation.Taken together, these patterns confirm that meta-learning is most stable within a mid-capacity regime. Models must be large enough to encode reusable structure, but not so large that their learning becomes erratic. These insights help contextualize downstream findings: the best generalization often arises from models that strike a balance between representational power and stable task-level adaptation.
@@
To probe how episodic adaptation reshapes the backbone’s feature geometry, we track the mean and standard deviation of the episodically adapted classifier head across training (Figure 4). Because the inner loop updates only this shallow head on frozen backbone features, its across-episode weight statistics act as a lightweight linear-probe proxy for class separability: under softmax on fixed features, class weight vectors tend to align with differences between class means, so greater dispersion (std) across head weights indicates larger between-class margins induced by the backbone, while transient spikes without sustained query gains suggest support overfit rather than stable generalization. We therefore relate inflections in mean/std to simultaneous changes in support/query accuracy to contextualize adaptation quality.The top panel shows the standard deviation of head weights. All models exhibit growth in weight variance, indicating increasing expressivity in the task-specific head. The small model diverges most sharply, with its weight variance surpassing all others after 2k steps. This suggests an over-specialization effect: the model learns to adapt aggressively to each task, potentially at the cost of stability.In the lower panel, the mean of the head weights remains near zero for most models, but the tiny model is an outlier. It accumulates a strong bias in one direction over training, indicating that its head converges toward a fixed mapping that is minimally updated across episodes. This aligns with earlier diagnostics showing that its gradient norms collapse early in training.These dynamics reinforce the idea that episodic MAML indeed induces a scale-sensitive tradeoff: in higher-capacity models, episodic gradients drive generalizable structure into the shared initialization; in lower-capacity models, this same pressure can cause drift or collapse.
@@
We interpret this behavior as a representational phase transition: the model initially fits the objective using diffuse, high-dimensional representations, which are later compressed into task-specialized, low-rank structures. The descent in PER lags behind the initial perplexity gains, and only after this drop does the second descent in Paloma begin. There is no strong evidence of a comparable phase transition in the vanilla models. While the large and medium variants show mild inflection points in loss and perplexity around step 3000, these are gradual and lack the coordinated sharpness seen in the MAML-trained models.This suggests that MAML’s bilevel updates and episodic task pressure may help reorganize the optimization landscape to favor discrete qualitative shifts in representation. As explored in Olsson et al. (2022); Wang et al. (2024); Hoogland et al. (2024), model training often proceeds in qualitatively distinct stages: from brute-force fitting, to intermediate rule memorization, to compressed algorithmic abstraction. The drop in PER may signal such a transition—from early diffuse representations to compressed heads tuned to solve the repeated structure of SMLMT episodes. This representational transition is also reflected in the model’s adaptation performance. Around the same step where PER and Paloma perplexity undergo a sharp drop (step 3000), both support and query accuracies rise abruptly (see Figure 3). Prior to this point, query accuracy remains relatively flat, indicating that the model struggles to generalize from support to query examples. But after the phase transition, the model rapidly learns to extrapolate, with query accuracy climbing from near random to over 0.5.
This synchrony across metrics provides compelling evidence of a coordinated phase shift in the model’s learning trajectory. When looking into more granular checkpoints (Figure 6), there is clearer evidence that the model transitions from an early stage where it relies on diffuse representations to a later stage where it reorganizes both its representations and update paths into a lower-dimensional, more modular form capable of few-shot generalization. That said, this phase behavior appears scale-sensitive as it is absent in smaller scales. This suggests that the capacity to reorganize may be gated by scale, and that below a certain threshold, the inductive pressure of MAML induces collapse rather than modularization.
@@ 7 Conclusion
We interleaved first–order MAML episodes with decoder pretraining and analyzed dynamics across four SLM scales. Under equal compute, the hybrid objective accelerates optimization but trades off perplexity at most scales; downstream, it brings modest average NER gains (+2–3 pp) at medium/large scales. Spectral logs expose a phase-like diversify–then–compress pattern that coincides with improving episodic query accuracy in the large model. Given our NER-only, single-seed scope, we present these as tools and observations rather than broad performance claims.However, while our evaluation focuses exclusively on named entity recognition, the underlying mechanism—episodic adaptation via SMLMT—is task-agnostic. In principle, the same hybrid objective could be applied to other sequence labeling tasks (e.g., part-of-speech tagging, syntactic chunking) or even structured prediction problems that admit few-shot formulations. Whether the phase transitions and rank-compression patterns we observe generalize to non-linguistic domains (e.g., code generation, mathematical reasoning) remains an open question. Future work should explore whether MAML’s inductive bias is inherently suited to token-level structure learning or whether it confers broader benefits across modalities and task families.Relatedly, other natural extensions suggest themselves. Future work should also include multi-seed and hyperparameter sweeps (inner LR, episode frequency), multilingual pretraining to test cross-script transfer, varying which layers adapt in the inner loop, and evaluation on non-NER tasks (e.g., classification, QA, reasoning), as the architectural design space is rather large. In terms of exploratory work, a natural next step is to learn whether the same phase transition re-emerges when the corpus is multilingual, which would clarify why cross-script transfer remains the weak point of the present models. Varying which backbone layers adapt, how many steps they receive and how frequently episodes are interleaved may unlock better compute–capability trade-offs. Finally, the clear correlation between the effective-rank collapse and downstream utility hints that spectral diagnostics might serve as a self-supervised early-stopping signal.
@@ Acknowledgments
This work was supported by a grant from the Accelerate Programme for Scientific Discovery, made possible by a donation from Schmidt Futures. David Demitri Africa is supported by the Cambridge Trust and the Jardine Foundation. Richard Diehl Martinez is supported by the Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation). Suchir Salhan is supported by Cambridge University Press & Assessment.